OpenStudy (varaug):

kelly has a $3,000 balance on her credit card with an 18% interest rate. if she makes no payments on her card and no late fees are charged, how long will it take for her debt to double if its compounded quarterly?

OpenStudy (varaug):

@LGND_88

OpenStudy (varaug):

halp pls @wanell

OpenStudy (varaug):

@3mar

OpenStudy (varaug):

@paki

OpenStudy (varaug):

i know that the formula for interest is \[A=P _{0}(1+\frac{ r }{ n })^{n*t}\]

OpenStudy (varaug):

but i don t know how to set it up

OpenStudy (varaug):

*compound interest i mean*

OpenStudy (varaug):

@Luigi0210 @sammixboo @Awolflover1

OpenStudy (varaug):

@MathSucks321

OpenStudy (mathsucks321):

I suck at these sorry T_T Plus I have to finish my work that I decided to start today! And it is due Friday.

OpenStudy (varaug):

i would probably be able to do this on my own but i have ADHD so i need music to focus......at the same time the music makes it difficult for me to think. (cuz i listen to hardstyle, dubstep, etc) so its a trade-off. do i want to be able to think and not focus, or focus and not be able to think. T_O

OpenStudy (varaug):

@dehydrated

OpenStudy (varaug):

@pooja195

OpenStudy (varaug):

@zepdrix @IrishBoy123

OpenStudy (varaug):

i need help ASAP

OpenStudy (varaug):

help me out @518nad please.....been waitin for over an hour

OpenStudy (518nad):

kelly has a $3,000 balance on her credit card with an 18% interest rate. if she makes no payments on her card and no late fees are charged, how long will it take for her debt to double if its compounded quarterly?

OpenStudy (518nad):

okay this is something financial people do and its a bit.. ambiguous, when they say 18% interest rate compounded quarterly, what they really mean is 18/4% interest rate is compouneded every single quarter

OpenStudy (varaug):

alright...

OpenStudy (518nad):

18/4 = 4.5%

OpenStudy (518nad):

so every quarter ur compounding 4.5%

OpenStudy (varaug):

k.

OpenStudy (varaug):

so 4.5% of 3000 is........

OpenStudy (varaug):

crap forgot the formula...

OpenStudy (518nad):

3000(1+0.045) ^---------- 4.5%

OpenStudy (varaug):

ok. hang on...

OpenStudy (518nad):

so thats what u have to pay after 1 quarter

OpenStudy (518nad):

3000(1+0.045)(1+0.045) after 2 quarters 3000(1+0.045)(1+0.045)(1+0.045) after 3 quarters 3000(1+0.045)^n after n quarters u want to know how many quarters it takes to get to double of 6000 3000(1+0.045)^n=6000, solve for n

OpenStudy (varaug):

k gimmie a sec.

OpenStudy (varaug):

srry got kicked

OpenStudy (518nad):

3000(1+0.045)^n=6000

OpenStudy (varaug):

ehhhhmmm.... i got 1.49577e118

OpenStudy (518nad):

3000/3000(1+0.045)^n=6000/3000 (1+0.045)^n=2 1.045^n=2 n log 1.045 = log 2 n=log2/log1.045

OpenStudy (518nad):

log2/log1.045=15.7473 15.7474 quarters, so thats almost 4 years

OpenStudy (varaug):

wow. thx

OpenStudy (varaug):

i got another one if ur up to it

OpenStudy (518nad):

ya if u really understand this question

OpenStudy (varaug):

What if kellys APR was 22% and she makes no payments on her card and no late fees were charged? (use the same information from #15 to find out how many years it would take her 3000 debt to double)

OpenStudy (518nad):

okay try to do this urself, show me ur work

OpenStudy (518nad):

the only difference here is the 18% is now 22%

OpenStudy (varaug):

ok, im not going to lie, i dont understand your descripion. can you use the equation tab to show me how to do this in formula form?

OpenStudy (518nad):

okay lets start with this

OpenStudy (518nad):

lets consider a simpler example 40% annual rate, how much percentange is this compounded quarterly?

OpenStudy (varaug):

10%

OpenStudy (518nad):

right

OpenStudy (518nad):

so now lets see whats going on here, that means every quarter we have a 10% increase in what we curently have

OpenStudy (varaug):

which is $30

OpenStudy (518nad):

First quarter we must pay a total of 3000(1+0.1)

OpenStudy (varaug):

or $300 my bad

OpenStudy (518nad):

its 3300 total u have to pay

OpenStudy (varaug):

ah

OpenStudy (518nad):

one step at a time, do u understand this: First quarter we must pay a total of 3000(1+0.1)

OpenStudy (518nad):

hey?

OpenStudy (518nad):

please be straight forward its okay to say u dont understand if u dont

OpenStudy (varaug):

srry replying to a message

OpenStudy (518nad):

okay can u pay attention for 5 mins

OpenStudy (518nad):

if u can bear with me for 5 mins ill teach u everything u need to know about compounding that ull be able to derive urself

OpenStudy (varaug):

i would probably be able to do this on my own but i have ADHD so i need music to focus......at the same time the music makes it difficult for me to think. (cuz i listen to hardstyle, dubstep, etc) so its a trade-off. do i want to be able to think and not focus, or focus and not be able to think. THANK YOU VERY MUCH!!!! >:(

OpenStudy (518nad):

okay lets forget about everything that tells u u cant do this right now and focus on the question im asking u its simple

OpenStudy (518nad):

Do you understand : First quarter we must pay a total of 3000(1+0.1)

OpenStudy (varaug):

yea

OpenStudy (varaug):

simple equation

OpenStudy (varaug):

so far at least...

OpenStudy (518nad):

do u understand this first quarter 3000+3% of 3000

OpenStudy (varaug):

yes

OpenStudy (518nad):

3000+10% of 3000 3000+0.1*3000, factor the 3000 3000(1+0.1)

OpenStudy (518nad):

this is task 1 to get into this simplified form

OpenStudy (varaug):

OOOOH. ok i get it. you got the 0.1 by moving the decimal 2 times to the left becaise of the 100% thing. ok cool.

OpenStudy (518nad):

yes

OpenStudy (518nad):

okay now are u sureee u get 3000(1+0.1)

OpenStudy (518nad):

other wise the rest will confuse u

OpenStudy (varaug):

yea

OpenStudy (varaug):

shoot

OpenStudy (518nad):

let me just explain just ot be sure again

OpenStudy (varaug):

T_T fine

OpenStudy (518nad):

3000(1+0.1) the (1+0.1) what is that doing to this 3000? its taking that 3000 and then adding 10% of that 3000 to itself

OpenStudy (varaug):

ok.

OpenStudy (518nad):

3000(1+0.1), so this is the current total to pay at first month, at the 2nd month u must tke this whole price and add 10% of to that new price

OpenStudy (518nad):

which is why u multiply by another (1+0.1) 3000(1+0.1)(1+0.1) price to pay for 2nd quarter

OpenStudy (518nad):

does that make sense?

OpenStudy (varaug):

yea. one ques, is this calculator-safe?

OpenStudy (varaug):

the equation imean

OpenStudy (518nad):

hmm what?

OpenStudy (varaug):

like, can you input this into a TI-nspire calculater and it will work?

OpenStudy (518nad):

3000(1+0.1)(1+0.1) =3000(1+0.1)^2 for 2 quarters 3000(1+0.1)^n for n quarters

OpenStudy (518nad):

ya um try to understand this logic

OpenStudy (518nad):

3000(1+0.1)^n P(1+r)^n thus u get this equation

OpenStudy (varaug):

oh ok so the formula would be 3000(1+0.22)^n?

OpenStudy (518nad):

:<

OpenStudy (518nad):

yes, but remember n is the number of quarters here

OpenStudy (varaug):

which is unknown atm

OpenStudy (518nad):

yes so solve for n, and divide by 4 to turn the quarters into years

OpenStudy (varaug):

or ots 15.7474

OpenStudy (varaug):

oh

OpenStudy (518nad):

thats for hte previous question not this one

OpenStudy (varaug):

yea i figured too late